OpenStudy (anonymous):

Find the product of (4x − 3y)2.

OpenStudy (unklerhaukus):

\[ (4x − 3y)^2=(4x − 3y)(4x − 3y)=4x(4x − 3y)-3y(4x − 3y) \]

OpenStudy (unklerhaukus):

\[=16x^2-12xy-12yx+9y^2\]\[=16x^2-24xy+9y^2\]